HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

HIGH RESOLUTION MOLECULAR SPECTROSCOPY IN ORION WITH THE JWST

\underline{O. BERNÉ}, Institut de recherche en astrophysique et planétologie (CNRS and Univ. Toulouse). Avenue du Colonel roche, 31028, Toulouse, France.)

As part of the PDRs4All observing program (pdrs4all.org ), we have observed the Orion star-forming region using the JWST in imaging and spectroscopic modes. With a resolution of \frac{ν}{Δν} ∼ 3000, the spectra of the Orion Bar and the protoplanetary disks we have obtained provide access to hundreds of lines and bands from numerous atomic ions, neutral atoms, small and large molecules. The molecular species include the ro-vibrational transitions of H_{2}, HD, CO, CH^{+}, OH, and the emission from vibrationally excited polycyclic aromatic hydrocarbons.

In a protoplanetary disk, we detected a series of strong lines that did not correspond to any known species in spectroscopic databases. Combining efforts from astronomers and spectroscopists (both lab and theory), we were able to identify the molecular species responsible for the observed lines, i.e., the Methyl Cation CH_{3}^{+}. CH_{3}^{+} had been postulated to play a pivotal role in interstellar chemistry since the ’80s but remained elusive. We found that CH_{3}^{+} results from the hot UV-driven chemistry.

Overall, these first results demonstrate the spectroscopic potential of the JWST but also the necessity to rely on state-of-the-art spectroscopic data to interpret observations.

Bibliography

O. Berné, E. Habart, E. Peeters and the PDRs4All team 2022, PDRs4All: A JWST Early Release Science Program on radiative feedback from massive stars, PASP 134, 1035

O. Berné, M. A. Martin-Drumel, I. Schroetter et al. 2023 Formation of the Methyl Cation by Photochemistry in a Protoplanetary Disk, Nature in press.

E. Habart, E. Peeters, O. Berné and the PDRs4All team 2023 NIR and MIR imaging view of the Orion Nebula, A&A in press.

R. Chown, A. Sidhu, E. Peeters and the PDRs4All team 2023 An embarrassment of riches: the Aromatic Infrared Bands in the Orion Bar, A&A in press.

p-number: p119  Submitted on Wed, 14 Jun 2023 10:16:33 +0200